OpenStudy (anonymous):

If the nth partial sum of a sequence an is given by n Σ 2k+4 k=1 then what is the 4th term of the sequence?

OpenStudy (anonymous):

4th term is 12. 4th partial sum is 36.

OpenStudy (anonymous):

\[\sum_{k = 1}^{4} 2k+4 =(2(1)+4)+(2(2)+4)+(2(3)+4)+(2(4)+4) = 36 \]